Test 2 Timelimit: 1 hour. 15 minutes. 1:08:11 remaining. A particular fruits weights are normally distributed, with a mean of 445 grams and a standard deviation of 40 grams. The heaviest 7% of fruits weigh more than how many grams? Glve your answer to the nearest gram. C 91(0/1) 2 (01) 95(a/1) Points possible: 1 This is attempt 1 of 98(01) Submit 911 (01) 912 (01 913(01) Grade: 0/13 Print Version
Answer #1

havisest 7% will fall above 93rd percentile for which critical value z =1.4758

therefore corresponding weight =mean +z*std deviation =445+1.4758*40=504 gm